# Explaining Large Language Models with gSMILE: Generative Statistical Model-Agnostic Interpretability with Local Explanations

### v2

@@ 4.3 Developing the Interpretable Surrogate Model
To interpret the influence of individual words on text generation, we map the output space of the generation model to a one-dimensional distance scale, where each point corresponds to a perturbed prompt [128]. This enables the training of a local surrogate model that relates changes in input structure to variations in output behavior.
In this weighted linear regression framework, the perturbation vectors (i.e., binary or one-hot encodings of included or excluded tokens) serve as independent variables. The response variable is the Wasserstein distance between the output of a perturbed prompt and that of the original prompt. Each perturbation is assigned a weight based on its semantic similarity to the original prompt [114, 129].
The loss function minimized during training is a weighted squared error:
$\text{Loss}(\phi_{\text{gen}},\delta_{\text{WMD}},p_{\text{org}})=\frac{1}{n}\sum_{i=1}^{n}\pi_{i}(p_{\text{org}},p_{\text{pert}}^{i})\cdot\left(W_{i}(t_{\text{org}}^{\prime},t_{\text{pert}}^{\prime i})-\hat{f}_{t}(p_{\text{pert}}^{i})\right)^{2}$ (16)
Here:
  • $\phi_{\text{gen}}$ is the text generation model,
  • $\delta_{\text{WMD}}$ denotes the function used to embed generated texts,
  • $W_{i}(t_{\text{org}}^{\prime},t_{\text{pert}}^{\prime i})$ is the Wasserstein distance between the generated outputs,
  • $\pi_{i}$ is the similarity-based kernel weight for the $i$-th perturbation,
  • and $\hat{f}_{t}(p_{\text{pert}}^{i})$ is the prediction from the linear surrogate model.
The surrogate model $\hat{f}_{t}$ is trained on the one-hot vector representation of the perturbation prompts $p_{\text{pert}}^{i,01}$, using the computed Wasserstein distances as targets [114].
@@ 4.4 Evaluation Metrics
In our investigation, we adopt a suite of evaluation metrics inspired by foundational work provided by Google [130], emphasising the multifaceted nature of assessing explainable models. This work highlights the significance of metrics such as accuracy, stability, fidelity and consistency as essential tools for a rigorous evaluation of model behaviour, particularly when comparing explainable models to traditional black-box models. The adoption of these metrics provides a structured methodology to dissect and understand model reliability in a more holistic manner [130].
Figure 6:
4.4.1 Accuracy
Accuracy measures how well the model’s output aligns with the expected results (ground truth). Specifically, we compare the model’s attribution scores (attention to text elements) against ground truth labels that identify the most relevant text elements for specific generated text segments [131].
@@ 4.4.1 Accuracy
4.4.2 Stability
Stability in an explainable model refers to its ability to provide consistent explanations even when minor changes are made to the input data [134]. In other words, small perturbations in the input should not significantly affect the model’s predictions or explanations. Stability is a key property for fostering trust in explainable artificial intelligence or XAI systems [61]. It reassures users that the model’s interpretations are not arbitrary or overly sensitive to irrelevant variations [10].
In the example text, the attention scores are visualised for two slightly different text prompts. The heatmaps show how the model assigns attention to each word. A stable model, as demonstrated in Fig. 6, assigns similar importance to the words make and rainy in both prompts, even though the token  was added in one case. This consistency is critical for ensuring the model behaves predictably across minor input variations [136]. Without such stability, the explanations generated by the model could lead to confusion or misinterpretation, especially in high-stakes scenarios such as medical diagnosis or financial decision-making [63].
We use the Jaccard index to quantify stability, which measures the similarity between two sets of model predictions. For two sets $A$ and $B$, the Jaccard index is calculated as follows in Eq. 17:
@@ 4.4.2 Stability
This metric reflects the proportion of shared elements between the sets, where values closer to 1 indicate greater stability and similarity in explanations [134]. A higher Jaccard index confirms the model’s ability to yield stable outputs, thus providing consistent and reliable interpretations [4]. By ensuring stability, we not only enhance the robustness of the model but also its usability and credibility in real-world applications [61].
4.4.3 Consistency
In evaluating the consistency of our model, we assess whether it produces stable and reliable outputs when the same input is provided multiple times. Consistency ensures that the model behaves predictably, reinforcing trust in its outputs and usability across diverse scenarios [63].
As shown in Fig. 6, when the input phrase “what is the meaning of life?” was fed into the model multiple times, the output remained stable across iterations. This consistency is crucial for applications where repeatable and reliable results are required. In our case, the model consistently assigns higher weights to the words “meaning” and “life”, demonstrating that it understands the importance of these words in generating the desired output [4]. Such behaviour underscores the model’s ability to focus on the semantic core of the input, a trait essential for maintaining coherence in practical use cases [10].
In practical applications—such as automated content generation, interactive systems, or scenario simulations—users depend on models to produce results that do not fluctuate unexpectedly. Any inconsistencies could undermine user trust and compromise the system’s overall effectiveness. Our model enhances user confidence by demonstrating consistent outputs and establishes a foundation for further optimisation and integration into more complex systems. This ability to maintain consistency across iterations directly contributes to the robustness and scalability of the model, making it well-suited for deployment in dynamic real-world environments [61].
4.4.4 Fidelity
Fidelity is a metric used to evaluate how closely an explainable model replicates the behavior of a black-box model. Specifically, the predictions of the black-box model denoted as $f(X_{i})$, are compared to those of the explainable model, $g(X_{i})$, for each input perturbation $i$. A variety of accuracy-based and error-based metrics are employed to assess this alignment, thereby quantifying how well the explainable model approximates the responses of the black-box model.
@@ 4.4.4 Fidelity
$R^{2}=1-\frac{\sum_{i=1}^{N_{p}}(f(X_{i})-g(X_{i}))^{2}}{\sum_{i=1}^{N_{p}}(f(X_{i})-\bar{f})^{2}}$ (18)
where $\bar{f}$ represents the mean of the black-box model’s predictions $f(X_{i})$. A value of $R^{2}$ close to 1 indicates a strong alignment between the models.
@@
$\text{MSE}=\frac{1}{N_{p}}\sum_{i=1}^{N_{p}}(f(X_{i})-g(X_{i}))^{2}$ (21)
$\text{MAE}=\frac{1}{N_{p}}\sum_{i=1}^{N_{p}}|f(X_{i})-g(X_{i})|$ (22)
For situations requiring sample-level importance, the Weighted Mean Squared Error (WMSE) and Weighted Mean Absolute Error (WMAE) are defined as [138]:
$\text{WMSE}=\frac{\sum_{i=1}^{N_{p}}w_{i}\cdot(f(X_{i})-g(X_{i}))^{2}}{\sum_{i=1}^{N_{p}}w_{i}}$ (23)
$\text{WMAE}=\frac{\sum_{i=1}^{N_{p}}w_{i}\cdot|f(X_{i})-g(X_{i})|}{\sum_{i=1}^{N_{p}}w_{i}}$ (24)
Here, $w_{i}$ denotes the sample weight associated with the $i$-th perturbation.
@@
$L_{1}=\frac{1}{N_{p}}\sum_{i=1}^{N_{p}}|f(X_{i})-g(X_{i})|$ (25)
$L_{2}=\frac{1}{N_{p}}\sum_{i=1}^{N_{p}}(f(X_{i})-g(X_{i}))^{2}$ (26)
These metrics, MSE, MAE, WMSE, WMAE, $L_{1}$, and $L_{2}$—together quantify how accurately the explainable model reproduces the black-box model’s predictions and reflect the risk of prediction discrepancies across the perturbation space.
We compute fidelity across various scenarios by comparing the explainable model’s predictions to those of the black-box text generation model. This includes analyzing how different input perturbations affect fidelity scores and comparing the performance of various distance metrics and surrogate models, such as linear regression and Bayesian ridge regression, in aligning with the black-box behavior.
@@ 5 Experimental Results
This section evaluates the proposed method’s ability to enhance explainability in instruction-based text generation. We analyse how the structure of textual prompts affects interpretability, how variations in input text influence the clarity of explanations, and how the proposed solution performs across diverse scenarios.
Experiments were conducted using a range of datasets and three state-of-the-art text generation models, OpenAI GPT [1], LLaMA [19], and Claude-AI [140] to ensure a comprehensive assessment.
@@ 5.1 Qualitative Results
The qualitative evaluation demonstrates the proposed framework’s interpretability using scenario-based testing and visualisations. We designed diverse testing scenarios by varying textual prompts, including changes in sentence complexity, rephrasing, and the inclusion of domain-specific terminology. These scenarios examine how the model interprets and generates outputs based on nuanced variations in textual inputs.
@@ 5.1.1 Input Prompt Explainability
The proposed framework applies scenario-based testing across multiple prompt variations to evaluate the interpretability of the model’s behaviour regarding different inputs. These variations include rephrased questions, descriptive modifiers, and changes in word choice. The system generates a text heatmap highlighting individual words’ contribution to the model’s output for each input, with colours indicating the importance scores.
Figure 7 shows how varying the input prompt’s phrasing or focus leads to different word importance patterns, demonstrating the model’s sensitivity to linguistic changes.
@@ 5.1.2 Reasoning Path and Token Attribution
Beyond token-level importance visualisations, the proposed framework provides structured reasoning paths that illustrate how the model logically connects input elements to produce its response. Figure 8 presents an example where the gSMILE method not only assigns attribution scores to individual tokens but also generates a reasoning flow diagram. This combined visualisation offers a holistic understanding of local word-level contributions and the logical structure of the model’s decision-making.
Such comprehensive visual explanations align with advanced techniques used by leading AI research organisations, including Anthropic and OpenAI, while providing greater flexibility across diverse tasks and models. As shown in Figure 8, gSMILE’s explainability aligns with or exceeds the interpretability achieved by leading industry methods [141].
Figure 8: Comparison between attribution visualisations by Anthropic’s Attribution Graphs and the gSMILE framework. The upper section shows the reasoning path identified by Anthropic’s method, while the lower section illustrates gSMILE’s token-level importance heatmap for the same prompt. [141]
@@ 5.1.3 Word-Level Attribution Analysis on MMLU Using gSMILE
To evaluate the interpretability of ChatGPT using the gSMILE framework, we used examples from the MMLU (Massive Multitask Language Understanding) dataset [142]. This benchmark contains four-choice multiple-choice questions covering various academic and professional subjects, including science, history, law, and the humanities. For this study, we accessed a pre-formatted version of the dataset available on Kaggle [143] to facilitate prompt construction and experimentation.
Figure 9 shows a set of attribution heatmaps that visualise which parts of the input prompt most influenced the model’s decision. Darker red tones indicate higher influence, while neutral or irrelevant tokens are shaded lighter. These visualisations demonstrate gSMILE’s ability to reveal the model’s internal reasoning by tracing the most impactful words in multi-choice question contexts.
Figure 9: Attribution heatmaps generated by gSMILE using ChatGPT (GPT-4) on selected samples from the MMLU dataset. Highlighted words represent those with the highest influence on the model’s output.
@@ 5.1.4 Optimised Instructions from Google Research
To further explore how instruction phrasing influences interpretability, we applied gSMILE to analyse two high-performing prompts introduced by Google Research through their OPRO framework [144]. These prompts were designed to enhance reasoning in LLMs and were explicitly optimised for zero-shot performance on mathematical tasks.
We first applied gSMILE to GPT-3.5 using 1024 perturbations on the instruction: “A small quantity of arithmetic and a logical approach will help us quickly solve this problem. ” As shown in Figure 10, the most influential words identified were “arithmetic” and “approach,” and “solution”, confirming their importance in shaping reasoning-oriented outputs.
We then extended the analysis to GPT-4 using the instruction: “Let’s combine our numerical command and clear thinking to quickly and accurately decipher the answer.” As illustrated in Figure 11, gSMILE revealed that “numerical,” “command,” and “answer” had the highest positive influence.
@@
Figure 11: Word-level importance in an instruction optimized by Google Research, analysed using gSMILE with GPT-4.

### v4

@@ 4.3 Developing the Interpretable Surrogate Model
To interpret the influence of individual tokens on the model’s output, we map the output space of the black-box text generation model $\pi^{(n)}$ to a one-dimensional distance scale, where each point corresponds to a perturbed prompt [118]. This mapping enables the training of a local surrogate model that relates structural changes in the input prompt to variations in output behaviour.
In this weighted linear regression framework, the independent variables are binary (or one-hot) perturbation vectors indicating which tokens from the original prompt $x$ are retained or altered in $\hat{x}_{j}$. The dependent variable is the output-level distribution shift $\Delta(x,\hat{x}_{j})$, defined in Eq. 3 as the Wasserstein distance between the output distributions for $x$ and $\hat{x}_{j}$. Each perturbation is assigned a similarity-based weight $w_{j}$ from Eq. 4, computed via the Gaussian kernel over the input-level distance $\delta_{x_{j}}$.
The surrogate model parameters are estimated by minimising the weighted squared error:
$\mathcal{L}(\pi^{(n)},\delta_{x_{j}},x)=\frac{1}{J}\sum_{j=1}^{J}w_{j}(x,\hat{x}_{j})\cdot\left(\Delta(x,\hat{x}_{j})-h_{\theta}(z_{j})\right)^{2}$ (13)
The fitted coefficients $\theta$ quantify the contribution of each token to the observed output shift, enabling interpretable attribution.
We evaluate the proposed attribution method using several state-of-the-art large language models: OpenAI’s gpt-3.5-turbo-instruct [54], Anthropic’s Claude 2.1 [29], and Meta’s LLaMA 3.1 Instruct Turbo (70B) [55].
@@ 4.4 Evaluation Metrics
In our investigation, we adopt a suite of evaluation metrics inspired by foundational work provided by Google [119], emphasising the multifaceted nature of assessing explainable models. This work highlights the significance of metrics such as ATT accuracy, ATT stability, ATT fidelity, ATT faithfulness, and ATT consistency as essential tools for a rigorous evaluation of model behaviour, particularly when comparing explainable models to traditional black-box models, as you can see in Fig. 6. The adoption of these metrics provides a structured methodology to dissect and understand model reliability in a more holistic manner [119].
4.4.1 Attribution Accuracy
Att Accuracy measures how well the model’s output aligns with the expected results (ground truth). Specifically, we compare the model’s attribution scores (attention to text elements) against ground truth labels that identify the most relevant text elements for specific generated text segments [120].
@@ 4.4.1 Attribution Accuracy
Figure 6: Given a prompt (e.g., “What is the meaning of life?”), We compare ground truth attributions with model-generated attributions for input tokens. Attribution quality is evaluated using four metrics: ATT Accuracy quantifies alignment of attributions with ground-truth (via AttAUC scores). ATT faithfulness measures whether attributions reflect the importance of tokens for the model’s response (via auROC). ATT consistency assesses whether results remain stable across different model runs or hyperparameters. ATT stability examines the sensitivity of attributions to small perturbations in the input prompt.
4.4.2 Attribution Faithfulness
ATT faithfulness refers to the extent to which an explanation accurately reflects a model’s reasoning process. A faithful attribution assigns high importance to input tokens that genuinely drive the model’s output, rather than highlighting spurious correlations or irrelevant features. This property is widely recognised as essential for trustworthy interpretability in natural language processing and other domains [123, 124].
In our study, we evaluate ATT faithfulness through correlation analysis between attribution-based metrics (ATT ACC, ATT F1, ATT AUROC) and externally reported benchmark accuracies of the models being assessed, including OpenAI’s gpt-3.5-turbo-instruct, Meta’s LLaMA 3.1 Instruct Turbo (70B), and Anthropic’s Claude 2.1 [125, 126, 127].
Formally, we compute the Pearson correlation coefficient:
$r=\frac{\sum_{i=1}^{n}\bigl{(}x_{i}-\bar{x}\bigr{)}\bigl{(}y_{i}-\bar{y}\bigr{)}}{\sqrt{\sum_{i=1}^{n}(x_{i}-\bar{x})^{2}}\,\sqrt{\sum_{i=1}^{n}(y_{i}-\bar{y})^{2}}},$ (14)
where $x_{i}$ are published benchmark ATT Accuracy values, $y_{i}$ are the corresponding attribution scores, and $\bar{x},\bar{y}$ denote their means. A strong positive $r$ indicates that attribution metrics faithfully track external measures of model competence.
Thus, ATT faithfulness in our framework is treated as an externally validated property: explanations are considered reliable to the extent that attribution scores correlate with established benchmarks, ensuring interpretability methods provide meaningful insights into large language models, particularly in high-stakes applications [128].
4.4.3 Attribution Stability
ATT stability in an explainable model refers to its ability to provide consistent explanations even when minor changes are made to the input data [129]. In other words, small perturbations in the input should not significantly affect the model’s predictions or explanations. ATT stability is a key property for fostering trust in explainable artificial intelligence or XAI systems [61]. It reassures users that the model’s interpretations are not arbitrary or overly sensitive to irrelevant variations [130].
In the example text, the attention scores are visualised for two slightly different text prompts. The heatmaps show how the model assigns attention to each word. A stable model, as demonstrated in Fig. 6, assigns similar importance to the words make and rainy in both prompts, even though the token  was added in one case. This ATT consistency is critical for ensuring the model behaves predictably across minor input variations [131]. Without such ATT stability, the explanations generated by the model could lead to confusion or misinterpretation, especially in high-stakes scenarios such as medical diagnosis or financial decision-making [64].
We use the Jaccard index to quantify ATT stability, which measures the similarity between two sets of model predictions. For two sets $A$ and $B$, the Jaccard index is calculated as follows in Eq. 15:
@@ 4.4.3 Attribution Stability
This metric reflects the proportion of shared elements between the sets, where values closer to 1 indicate greater ATT stability and similarity in explanations [129]. A higher Jaccard index confirms the model’s ability to yield stable outputs, thus providing consistent and reliable interpretations [8]. By ensuring ATT stability, we not only enhance the robustness of the model but also its usability and credibility in real-world applications [61].
4.4.4 Attribution consistency
In evaluating the ATT consistency of our model, we assess whether it produces stable and reliable outputs when the same input is provided multiple times. ATT consistency ensures that the model behaves predictably, reinforcing trust in its outputs and usability across diverse scenarios [64].
As shown in Fig. 6, when the input phrase “what is the meaning of life?” was fed into the model multiple times, the output remained stable across iterations. This ATT consistency is crucial for applications where repeatable and reliable results are required. In our case, the model consistently assigns higher weights to the words “meaning” and “life”, demonstrating that it understands the importance of these words in generating the desired output [8]. Such behaviour underscores the model’s ability to focus on the semantic core of the input, a trait essential for maintaining coherence in practical use cases [130].
In practical applications, such as automated content generation, interactive systems, or scenario simulations, users depend on models to produce results that do not fluctuate unexpectedly. Any inconsistencies could undermine user trust and compromise the system’s overall effectiveness. Our model enhances user confidence by demonstrating consistent outputs and establishes a foundation for further optimisation and integration into more complex systems. This ability to maintain ATT consistency across iterations directly contributes to the robustness and scalability of the model, making it well-suited for deployment in dynamic real-world environments [61].
4.4.5 Attribution Fidelity
ATT fidelity is a metric used to evaluate how closely an explainable model replicates the behaviour of a black-box model. Specifically, the predictions of the black-box model, denoted as $f(X_{j})$, are compared to those of the explainable model, $g(X_{j})$, for each input perturbation $j$. A variety of accuracy-based and error-based metrics are employed to assess this alignment, thereby quantifying how well the explainable model approximates the responses of the black-box model.
@@ 4.4.5 Attribution Fidelity
$R^{2}=1-\frac{\sum_{j=1}^{N_{p}}(f(X_{j})-g(X_{j}))^{2}}{\sum_{j=1}^{N_{p}}(f(X_{j})-\bar{f})^{2}}$ (16)
where $\bar{f}$ represents the mean of the black-box model’s predictions $f(X_{j})$. A value of $R^{2}$ close to 1 indicates a strong alignment between the models.
@@
$\text{WMSE}=\frac{\sum_{j=1}^{N_{p}}w_{j}\cdot(f(X_{j})-g(X_{j}))^{2}}{\sum_{j=1}^{N_{p}}w_{j}}$ (19)
$\text{WMAE}=\frac{\sum_{j=1}^{N_{p}}w_{j}\cdot|f(X_{j})-g(X_{j})|}{\sum_{j=1}^{N_{p}}w_{j}}$ (20)
Here, $w_{j}$ denotes the sample weight associated with the $j$-th perturbation.
@@
$L_{1}=\frac{1}{N_{p}}\sum_{j=1}^{N_{p}}|f(X_{j})-g(X_{j})|$ (21)
$L_{2}=\frac{1}{N_{p}}\sum_{j=1}^{N_{p}}(f(X_{j})-g(X_{j}))^{2}$ (22)
These metrics, MSE, MAE, WMSE, WMAE, $L_{1}$, and $L_{2}$, together quantify how accurately the explainable model reproduces the black-box model’s predictions and reflect the risk of prediction discrepancies across the perturbation space.
We compute ATT fidelity across various scenarios by comparing the explainable model’s predictions to those of the black-box text generation model. This includes analysing how different input perturbations affect ATT fidelity scores and comparing the performance of various distance metrics and surrogate models, such as linear regression and Bayesian ridge regression, in aligning with the black-box behaviour.
The overall workflow for computing ATT fidelity is shown in Figure 7. As depicted, input prompts are perturbed, and both the Global Model and the Local Surrogate Model produce similarity signals that are subsequently compared to evaluate ATT fidelity.
Figure 7: ATT fidelity evaluation framework. The input prompt is perturbed, and both a Global Model (e.g., ChatGPT) and a Local Surrogate Model (SMILE) generate similarity signals. Their comparison quantifies how closely the surrogate approximates the black-box model’s behaviour.
@@ 5 Experimental Results
This section evaluates the proposed method’s ability to improve explainability in instruction-based text generation, focusing on both performance and its implications for model transparency and usability. We analyse how prompt structure influences interpretability, how phrasing variations affect the clarity of explanations, and how the method performs across diverse scenarios.
For researchers, the analysis highlights linguistic factors that most affect interpretability; for practitioners, it offers actionable insights for prompt engineering and model debugging. Experiments were conducted using a range of datasets and three state-of-the-art text generation models, OpenAI’s gpt-3.5-turbo-instruct, Meta’s LLaMA 3.1 Instruct Turbo (70B), and Anthropic’s Claude 2.1, ensuring a robust and generalisable evaluation across architectures.
@@ 5.1 Qualitative Results
We assess interpretability through controlled variations of input prompts, including changes in sentence complexity, rephrasing, and the use of domain-specific terminology. This approach allows us to examine whether attribution patterns remain stable and meaningful under different linguistic conditions, a key requirement for trustworthy explainability tools. Observing these changes provides practical guidance for designing effective prompts and identifying potential reliability issues before deployment.
@@ 5.1.1 Input Prompt Explainability
The framework applies multiple prompt variations, such as rewording questions, adding descriptive modifiers, or altering word choice. It generates heatmaps showing the contribution of each token to the model’s output, with colour intensity indicating importance scores.
Figure 8 illustrates how small changes in prompt phrasing or focus can lead to noticeably different attribution patterns when using OpenAI’s gpt-3.5-turbo-instruct. These differences are not merely cosmetic: they reveal how the model’s attention can shift even for semantically similar inputs. Understanding such sensitivity helps practitioners design prompts that minimise unwanted variation, while also highlighting cases where model behaviour may be unstable, valuable both for prompt-engineering strategies and for reliability assessments in real-world applications.
@@ 5.1.2 Reasoning Path and Token Attribution
Beyond token-level heatmaps, the proposed framework generates structured reasoning paths that show how the model connects input elements to form its response. As illustrated in Figure 9, gSMILE not only assigns attribution scores to individual tokens but also produces a reasoning flow diagram, offering a unified view of both local word-level contributions and the broader logical structure behind the output.
This dual-layer explanation bridges the gap between fine-grained token attribution and higher-level reasoning, enabling users to trace how specific inputs influence each decision step. The approach aligns with advanced methods used by leading AI research organisations, such as Anthropic’s Attribution Graphs, while extending flexibility across diverse tasks and models. In the example shown, gSMILE’s interpretability matches or surpasses industry standards, providing a richer and more adaptable explainability toolkit for practical deployment [12].
Figure 9: Comparison between attribution visualisations by Anthropic’s Attribution Graphs (top) and the gSMILE framework (bottom) using OpenAI’s gpt-3.5-turbo-instruct. The former shows the reasoning path, while the latter presents token-level importance for the same prompt. [12]
@@ 5.1.3 Word-Level Attribution Analysis on MMLU Using gSMILE
We assessed the interpretability of OpenAI’s gpt-3.5-turbo-instruct with the gSMILE framework on the Massive Multitask Language Understanding (MMLU) benchmark [135], which contains four-choice questions across a broad range of academic and professional domains. A pre-formatted version from Kaggle [136] was used to streamline prompt creation and experimentation.
Figure 10 presents attribution heatmaps that highlight which tokens most strongly influenced the model’s choice. Darker red tones correspond to higher importance, while lighter shades indicate weaker influence. By surfacing these token-level attributions, gSMILE exposes the model’s internal focus when answering complex, domain-specific questions.
Such analysis is valuable for diagnosing whether the model bases its answers on semantically relevant cues or potentially spurious correlations. For example, consistent highlighting of key domain terms across questions suggests stable reasoning. In contrast, high attribution to irrelevant tokens may indicate susceptibility to misleading prompt elements, critical information for practitioners deploying LLMs in knowledge-intensive tasks.
Figure 10: Attribution heatmaps generated by gSMILE using OpenAI’s gpt-3.5-turbo-instruct on selected MMLU samples. Highlighted tokens are those with the greatest influence on the model’s prediction.
@@ 5.1.4 Optimised Instructions from Google Research
To examine how instruction phrasing affects interpretability, we applied gSMILE to two high-performing prompts from Google Research’s OPRO framework [137], which were optimised for zero-shot mathematical reasoning in LLMs.
For GPT-3.5, we analysed the instruction: “A small quantity of arithmetic and a logical approach will help us quickly solve this problem.” using 1024 perturbations. As shown in Figure 11, the tokens “arithmetic”, “approach”, and “solution” received the highest importance scores, indicating their role in directing the model toward structured problem-solving.
For GPT-4, we analysed the prompt: “Let’s combine our numerical command and clear thinking to quickly and accurately decipher the answer.” Figure 12 shows that “numerical”, “command”, and “answer” were most influential, suggesting that technical and goal-oriented terms firmly steer reasoning in high-performing instructions.
This Comparison highlights how effective prompt design amplifies semantically critical terms, improving reasoning reliability, a valuable insight for crafting robust instructions across different LLM architectures.
@@
Figure 12: Word-level importance in an instruction optimised by Google Research, analysed using gSMILE with GPT-4.